can someone please show me how to do these fractions? They are just plain adding or subtracting.I need to know how step by step how you work them please.
1) 1/4 + 2 7/8
2) 12 1/16 + 2 3/32
3) 4 2/3 - 1 1/4
4) 17 - 15/16
If you could explain how you do each step that would be so helpful to me.
Thank you so much in advance for your time.

Answer by Mathtut(3670)

your can change the mixed fractions to improper fractions but in order to add or subtract we need to make sure we have like terms(in other words the same number in the denominators or we can add the whole numbers together and then the fractional parts together. Lets try the improper fraction method.
1) 1/4 + 2 7/8
2) 12 1/16 + 2 3/32
3) 4 2/3 - 1 1/4
4) 17 - 15/16
:
:
1)2 7/8 = we know that 2 is 16/8 and then you add 7/8 so you have 23/8
:
so now we have (1/4)+(23/8). From this point we have to determine the Lowest COMMON denominator. since 8 is divisible by 4 and 8 , then 8 is our LCD. we need only change the first term divide 8 by 4 and then multiply the numberator in (1/4) by that answer so 8/4=2...1*2=2 so (1/4)=2/8. Now our problem looks like this (2/8)+(23/8)=25/8. Now we can leave this in form of an improper fraction or convert to a mixed fraction 25/8=3 1/8
:
2)in problem 2. I will show we have two options. We can either add the whole numbers together and then the fractions or go through the same process we went through in the last problem. lets try the whole number method
:
12 1/16 + 2 3/32 ..so 12 +2 =14. What we have left is (1/16)+(3/32). the LCD is 32, so we must convert the fraction 1/16 to a denominator of 32. 32/16=2. Multiply 2 by 1=2 so (1/16)=(2/32). So the fractional part now looks like this
:
(2/32)+(3/32)=5/32. So our answer is 14 5/32.
:
YOU TRY 3 and 4